OpenStudy (anonymous):

can someone help me put f(x)=2(x-3)^2+11 in standard form?

OpenStudy (anonymous):

First expand (x-3)^2 what will the equation look like?

OpenStudy (anonymous):

honestly i have no idea, can you help me understand it please? @nelsonjedi

OpenStudy (anonymous):

(x-3)^2 = (x-3) x (x-3) = what?

OpenStudy (anonymous):

seems like its an easy problem but im still so confused. @nelsonjedi

OpenStudy (anonymous):

Multiply (x-3) and (x-3).... x times x = ? , x times -3 = ?, -3 times x= ?, -3 times -3 = ?

OpenStudy (anonymous):

x2-6x+9 @nelsonjedi

OpenStudy (anonymous):

Very good now substitute that back into the problem for (x-3)^2 what do you have?

OpenStudy (anonymous):

f(x)=x2-6x+9+11 ? sorry if it's not right. @nelsonjedi

OpenStudy (anonymous):

Never say your sorry for trying you are close but go back and look at the original equation..what was in front of *x-3)^2 ??

OpenStudy (anonymous):

f(x)= 2x^2-6x+9+11 but then you add 9 and 11 right? so f(x)= 2x^2-6x+20 @nelsonjedi

OpenStudy (anonymous):

Close but it is 2(x^2-6x+9)+11 so everything in the parenthesis must be multiplied by 2.

OpenStudy (anonymous):

2x^2-12x+18+11 so 2x^2-12x+29 @nelsonjedi

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

thank you so much! (: @nelsonjedi

OpenStudy (anonymous):

You are welcome